5. Let p be the proportion of the population who respond yes.

a) In a survey of 700 US adults, 118 felt that alcoholism was a reasonable excuse for criminal conduct. Use the information to find p-hat & q-hat.

b) Find the maximum error of estimate for p, with a 93% confidence level (assume normally distributed).

c) Find the confidence interval with a 93% confidence level.
Answer #1

here x=118 n=700

p^=x/n=118/700=0.1686

q^=1-p^=1-0.1686=0.8314

Ansb:

ME=maximum error of estimate for p

AT 93 % CI the value of z is 1.8119

ME=Z*sqrt(p^*q^/n)

=1.8119*sqrt(0.1686*0.8314/700)

=0.0256

Ansc:

confidence interval with a 93% confidence level. is

(p^-ME,p^+ME)

(0.1686-0.0256,0.1686+0.0256)

(0.143,0.194)